OpenStudy (anonymous):

Given that f(x)=x^2+5 and g(x)=2x+7, calculate f∘g(x)= , g∘f(x)= , f∘f(x)= , g∘g(x)

OpenStudy (anonymous):

wow which one would you like to do first?

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(2x+7)=(2x+7)^2+5\] then algebra.. do you understand the steps?

OpenStudy (anonymous):

satellite what do you mean by algebra isnt that all to it

OpenStudy (anonymous):

i can write out the steps in english if you like?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh i mean probably you would write \[(2x+7)^2+5=4 x^2+28 x+54\]

OpenStudy (anonymous):

OK we want for example \[g\circ f(x)\] step one, get rid of the circle and write what this really means \[g(f(x))\]

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

step two, \[f(x)\] can be any function but now replace it by the specific one you are given. in this case it is \[f(x)=x^2+5\] so \[g(f(x))=g(x^2+5)\]

OpenStudy (anonymous):

then were you see an "x" in \[g(x)\] replace it by \[x^2+5\] using parentheses. in this case you get \[g(x^2+5)=(x^2+5)+7\]

OpenStudy (anonymous):

typo there, should have been \[g(x^2+5)=2(x^2+5)+7\]

OpenStudy (anonymous):

and finally algebra to clean it up

OpenStudy (anonymous):

so is it 2x^2 + 10 + 7

OpenStudy (anonymous):

well i would say \[2x^2+17\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

where it says g(g) and f(f)...thats strange. ive never seen that before. can you demonstrate it please

OpenStudy (anonymous):

it is exactly the same as before. \[f\circ f(x)=f(f(x))\]

OpenStudy (anonymous):

so again start from the inside, and work out \[f(f(x))=f(x^2+5)=(x^2+5)^2+5\]

OpenStudy (anonymous):

the reason this is confusing is because there are so many "x"s ignore them (mentally) if you can. for example the function \[g(x_=2x+7\] don't think "g of x equal two x plus seven" think "g means first double, then add seven"

OpenStudy (anonymous):

the function is the action, has nothing to do with the variable. so there is no difference between \[g(x)=2x+7\] and \[g(z)=2z+7\] and \[g(\heartsuit)=2\heartsuit+ 7\]